OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (\frac{ (1+x)^\frac{ 1 }{ x } }{ e })^\frac{ 1 }{ x }\]

OpenStudy (anonymous):

@AravindG Mr. Admin please give me solution not *answer* !!

OpenStudy (anonymous):

@AravindG ?

OpenStudy (anonymous):

@hartnn ?? @ganeshie8 ??

OpenStudy (anonymous):

@zepdrix ?? @Hero ??

OpenStudy (anonymous):

@AravindG ?

OpenStudy (anonymous):

@thomaster ?

OpenStudy (anonymous):

@beccaboo333 ?? @Luigi0210 ??

hartnn (hartnn):

let that limit be = L then take log on both sides and simplify the right side using log rules.

OpenStudy (anonymous):

u mean like log L=\[\frac{ 1 }{ x }(\log \lim_{x \rightarrow 0} (1+x)^\frac{ 1 }{ x }-\log \lim_{x \rightarrow 0} x)\]

OpenStudy (anonymous):

@hartnn ?

OpenStudy (anonymous):

@hartnn ??

OpenStudy (anonymous):

@Hero ?? @ganeshie8 ??

hartnn (hartnn):

sorry for late replies, my internet connection is very slow, don't you have a standard result for limit x \(\to 0 \) of log (1+x)^(1/x) ? if not, which standard limit formulas can you use ?

OpenStudy (anonymous):

the simplification i made isnt correct?

OpenStudy (anonymous):

can u simplify please!

OpenStudy (anonymous):

@hartnn ?

hartnn (hartnn):

\(\large \ln L = \lim \limits_{x\to0} (1/x) [\ln (1+x)^{(1/x)}-\ln e]\)

OpenStudy (anonymous):

lne=1

OpenStudy (anonymous):

@hartnn ?? now?

OpenStudy (anonymous):

@hartnn ?

OpenStudy (anonymous):

@hartnn ?